Klumper Corporation is a diversified manufacturer of industrial goods. The companys activity-based costing system contains the following six activity cost pcols and activity rates: Activity Cost Pool Supporting direct labor Activity Rates $ 9.00 per direct labor-hour $4.00 per machine-hour $45.00 per setup $ 160.00 per order $ 115.00 per shipment Machine setups Production orders Shipments Product sustaining 875.00 per product Activity data have been supplied for the following two products: Total Expected K425 M67 Number of units produced per year Direct labor-hours 1,150 3,000 200 2,000 50 40 Machine setups Production orders Shipments Product sustaining 14 Required Determine the total overhead cost that would be assigned to each of the products listed above in the activity-based costing system. Cost Pool K425 M67 Supporting direct labor Machine processing Machine setups Production orders Product sustaining
Answer #1

Answer:

Activity Cost Pool Calculation K425 Calculation M67
Supporting Direct Labor =1150*9 $10,350 =50*9 $450
Machine processing =3000*4 $12,000 =40*4 $160
Machine setups =7*45 $315 =3*45 $135
Production orders =7*160 $1,120 =3*160 $480
Shipments =14*115 $1,610 =3*115 $345
Product Sustaining =3*875 $2,625 =3*875 $2,625
Total Overhead Cost $28,020 $4,195